Solve the quadratic equation 5x2 - 125 = 0.
A. -5
B. 5
C. ±5
D. ±25

Answer by stanbon(75887)

Solve the quadratic equation 5x2 - 125 = 0.
Factor:
5(x^2-25) = 0
x^2-25 = 0
Factor:
(x-5)(x+5) = 0
----
And: x = 5 or x = -5
==========================
Cheers,
Stan H.
----------------------
A. -5
B. 5
C. ±5
D. ±25